Ottawa-Carleton Insitute of Mathematics and Statistics Workshop in Algebra (Lie and Finite-dimensional Algebras) September 27 and 28, 1996

Preliminary Program

Friday, September 27, Carleton University, Herzberg Laboratories, Room 4351. The afternoon will be devoted to associative algebras and algebraic groups. The speakers will be: 13:30 - B. Parshall (University of Virginia) 15:00 - E. Cline (University of Oklahoma) 16:30 - I.Ágoston (Hungarian Academy of Sciences /Carleton University)

Saturday, September 28, University of Ottawa, Department of Mathematics, 585 King Edward, Room B 005. The Saturday will be devoted to Lie algebras. The speakers will be: 9:00 - S. Berman (University of Saskatchewan) 10:30 - A. Pianzola (University of Alberta) 12:00 - G. Benkart (University of Wisconsin/ Institute for Advanced Study, Princeton)

14:00 - S. Futorny (University of Kiev/ Carleton University/University of Ottawa) 15:30 - K. Nishiyama (Kyoto University) 17:00 - O. Smirnov (Academy of Science, Novosibirsk/University of Ottawa)

There will be provision for further talks both on Friday and Saturday. The final program will be arranged on the spot. All participants will be invited to lunch on Saturday. For further information please contact Vlasta Dlab (vdlab@math.carleton.ca) or Erhard Neher (neher@uottawa.ca).